A theatre is holding an annual spring performance.Adults pay $7.50 and children pay $3.75.If there were 68 adult tickets sold,what is the minimum number of childrens tickets that need to be sold,in order for revenues to exceed $600?
The options include : 136,24,90,and 25

Answer by checkley77(12844)

68*7.50=$510 from sales of adult tickets.
$90 worth of children's tickets need to be sold.
90/3.75=24 children's tickets need to be sold to make $600.
To exceed $600 1 more children's ticket needs to be sold (24+1=25 ans.)